The speed of a passenger train is 18 mph faster than the speed od a freight train. The passenger train travels 250 miles in the same time it takes the freight train to travel 160 miles. Find the speed of each train.
What is the speed of passenger train?
What is the speed of freight train?

Answer by mananth(16946)

Let speed of freight train be x mph - 160 miles
passanger train speed = x+18 --------250
time is same for both
t= d/r
250/(x+18) = 160/x
250x= 160(x+18)
250x= 160x+ 2880
250x-160x=2880
90x= 2880
/90
x=32 mph
the speed of passenger train =32mph
the speed of freight train = 32+18 = 50 mph